OpenStudy (anonymous):

Help, I want to understand how you do this (not just the answer) and what formula(s) I have to use. 1.00ml of a 3.50x10^-4 M solution of oleic acid is diluted with 9.00mL of petroleum ether, forming solution A. Then 2.00mL of solution A is diluted with 8.00mL of petroleum ether, forming solution B. What is the concentration of solution B?

OpenStudy (aaronq):

You need to understand that you're diluting the sample. Use: \(M_1V_1=M_2V_2\) for example, you have 1 mL of \(3.5*10^{-4}\) M oleic acid and you add 9 mL of pet. ether new volume: 1 mL + 9mL = 10mL so, \(M_1V_1=M_2V_2 \rightarrow (3.5*10^{-4})*1 mL= M_2*(10mL) \rightarrow M_2=3.5*10^{-5}M\) so the concentration of solution A is \(3.5*10^{-5}M\). Apply the same concept for solution B

OpenStudy (anonymous):

But would you use the concentration of solution A or the concentration you were originally given in the problem?

OpenStudy (anonymous):

Nevermind the question says 2.00 ml of solution A so you'd use 3.5 * 10^-5

OpenStudy (anonymous):

Thank you so much!

OpenStudy (aaronq):

yep, use 2 mL of the new solution. no problem !